# From Path Coefficients to Targeted Estimands: A Comparison of Structural Equation Models (SEM) and Targeted Maximum Likelihood Estimation (TMLE)

### v1

@@ 1 Introduction
Structural Equation Modeling [36, SEM;] is a widely used framework in applied statistics, social sciences, psychology, and related fields [52, 62, 48, e.g.,], allowing researchers to model complex interrelationships among observed and latent variables. As a multivariate technique that often incorporates path analysis, SEM also provides an approach to capture mediational mechanisms through direct and indirect effects [39, 46]. In addition to SEM’s adaptability and flexibility, many software implementations are available (e.g., lavaan [59] in R; Stata; and Mplus), which offer further applicability for empirical research.
Numerous research questions in sociology and other fields such as psychology and education could be equivalently answered by estimating the path coefficients within the SEM framework. Under the structural equations and the distributional assumptions on the disturbance terms, parameters are typically estimated by maximizing the likelihood function induced by the model. The subsequent inferences are heavily based on likelihood-based tests or non-parametric approaches such as bootstrap [39, 36]. However, as the true data-generating mechanism typically remains unknown, it is not guaranteed that the specified likelihood encodes it correctly [39]. In the presence of model misspecification, the efficiency and consistency of the estimator and the validity of the corresponding statistical inference are no longer granted, resulting in unreliable estimates and inferences[49, 67].
@@
Notably, path coefficients and path diagrams (e.g., Figure 1) in SEM can sometimes represent causal effects that are defined under the Neyman–Rubin potential outcomes framework, linking causal inference to parameter estimation [37, 55, e.g.,]. However, the equivalence between path coefficients and causal effects essentially depends on the correct model specification. To address the limitations of linearity and reduce the risk of model misspecification, nonparametric structural equation models [56, NPSEM;] use nonparametric equations to represent relationships between variables, reformulating path diagrams as Directed Acyclic Graphs (DAGs; Figure 2). Using NPSEM and DAG, causal effects are identified as functionals of the joint distribution rather than path coefficients. Built on NPSEM, Targeted Maximum Likelihood (Loss) Estimation (TMLE), introduced by van der Laan and Rubin [64], provides a modern, semiparametric, machine learning-based framework for causal inference that is doubly robust and efficient under certain conditions.
@@
Upon identifying a causal effect from NPSEM, TMLE then involves two steps: an initial data-adaptive estimation followed by an adjustment that targets the estimation equation of the efficient influence function (EIF) [44]. TMLE is often coupled with the Super Learner algorithm [65]—an ensemble method that combines multiple learners to enhance predictions. Under regularity conditions, TMLE achieves asymptotic efficiency and hence provides a valid uncertainty measurement through the influence function [44, 64], offering computational advantages over bootstrap methods while preserving interpretability.
@@ 2.1 Structural Equation Modeling
[39] offers a concise review of SEM, from which the following overview is primarily adapted, with minor modifications for notation and clarity. The standard SEM consists of two parts: the measurement submodel and the structural submodel. In general, the structural equation is defined as
@@
Under regularity conditions, the properties of MLE ensure the asymptotic distribution of $\sqrt{n}(\hat{\boldsymbol{\theta}}_{\text{ML}}-\boldsymbol{\theta})$ is multivariate normal with mean $\boldsymbol{0}$ and covariance matrix given by the inverse of the Fisher information matrix, enabling statistical inference such as Wald tests for parameter significance and likelihood ratio tests for model fit. It is possible that the desired estimation involves some nonlinear transformations of the path coefficients. In such a case, the uncertainty is usually captured by the Delta method or non-parametric techniques like bootstrap [37].
@@ 2.2.1 An Overview of Targeted Learning
In contrast to parametric SEM, Targeted Maximum Likelihood(Loss) Estimation is a modern, semi-parametric estimation framework developed to provide an efficient and robust estimation of causal effects and other statistical parameters in complex data settings. An NPSEM is generally specified first for describing the causal relations and identifying the causal parameters in terms of the joint distribution. Then, the TMLE involves two main steps: an initial estimation of outcome functions using any suitable method, and a targeting step that updates these estimates to improve the estimation of the parameter of interest. This targeting is done through a fluctuation submodel and solving the estimation equation of the efficient influence function (EIF), which plays a key role in ensuring local efficiency. The resulting TMLE estimator is asymptotically linear and normally distributed under mild regularity conditions, allowing for valid inference via standard error estimation and confidence intervals. Following [45], we assume that a semi-parametric statistical model has been identified $\mathcal{M}$ with a true but unknown distribution $F_{0}\in\mathcal{M}$. The interested parameter is defined as $\Psi(F_{0}),\Psi(\cdot):\mathcal{M}\rightarrow\mathbb{R}$, and with $O_{1},O_{2},\cdots,O_{n}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}F_{0}$ observations from $F_{0}$. We further need that $\Psi(F_{0})=\Psi(Q_{0})$ depends on only $Q_{0}=Q(F_{0})$, a part of $F_{0}$, and $g_{0}$ is a nuisance parameter from some orthogonal factorization.
@@
which also yields a valid confidence set. This framework has been adapted to many estimands that admit the asymptotic linear estimation (e.g. [45, 47, 61]).
@@ 2.2.2 The Super Learner
As mentioned, the efficiency and consistency of TMLE depend heavily on the initial estimates of $Q_{n}^{0}$ and $g_{n}$. Though machine learning algorithms could be applied instead of the parametric regression, we by no means should rely on a single algorithm, as none is optimal for all datasets. It is worth mentioning that the estimation of $Q_{n}^{0}$ and $g_{n}$ can be done non-parametrically via the Super Learner [63], an ensemble learning technique that integrates modern machine learning algorithms. Consider a dataset $\{(X_{i},Y_{i})\}_{i=1}^{n}$, where $X_{i}\in\mathbb{R}^{p}$ are features and $Y_{i}\in\mathbb{R}$ are outcomes, jointly drawn from an unknown distribution $F_{0}\in\mathcal{M}$. Recall that the initial estimates in TMLE are equivalent to constructing a function $f(X)$ that minimizes the expected risk
@@
Figure 3: Illustration of the Super Learner algorithm flow. Adapted from [63].
Numerous modern machine learning algorithms are included in the current SuperLearner implementation in R, including Random Forest, Support Vector Machine, eXtreme Gradient Boosting, and Artificial Neural Network, which enhances its versatility and applicability. Notably, the choices of base learners and the meta learner play a key role in the final estimation, and have to be considered based on the specific tasks involved [57].
@@ 3 SEM and TMLE approach to causal inference
When considering SEM for causal inference, the commonly considered parameters are the Average Treatment Effect (ATE), Conditional Average Treatment Effect (CATE), and mediational effects [46, 37]. In this section, we consider these causal effects and show that both SEM and TMLE are applicable for estimation.
@@ 3.1 On Average Treatment Effects
Three essential causal assumptions for efficient identification and estimation of the causal effect from accessible data are (i) Stable Unit Treatment Value Assumption (SUTVA) or the consistency assumption: $Y=A\cdot Y(1)+(1-A)\cdot Y(0)$; (ii) no unmeasured confounders: $\mathbb{P}(A=1|Y(1),Y(0),\boldsymbol{X})=\mathbb{P}(A=1|\boldsymbol{X})$; and (iii) positivity: $\exists\epsilon>0\text{such that }\mathbb{P}(A=1|\boldsymbol{X}=\boldsymbol{x})>\epsilon$ for almost all $\boldsymbol{x}\in\mathcal{X}$, without which the ATE might not be identifiable [58].
@@ 3.1.1 The SEM approach to ATE
Suppose that the data-generating mechanism is encoded by the structural linear causal models [55]
@@
where $\ell$ denotes the log-likelihood function with $\phi(\mu,\tau^{2})$ denotes the normal density with mean $\mu$ and variance $\tau^{2}$, and $O$ represents the observed data, or as part of a generalized SEM with the logit link. It is unbiased, consistent, efficient, and asymptotically normal under correct model specification [66]. The statistical inference regarding the ATE $H_{0}:\gamma=0$, in a SEM context, is usually done via the Wald test and the likelihood ratio test.
@@
If $X_{k}=x_{k}$ is constant in $B$, the CATE simplifies to $\gamma+\tau x_{k}$. Estimation follows similarly, with asymptotic normality preserved under linear transformations, and the inference remains valid up to a minor modification [66].
@@ 3.1.2 The targeted learning approach to ATE
However, the identification of causal effects as path coefficients and the consistency of the estimators rely on correct functional form specification, and misspecification can introduce bias[39, 67]. In contrast to parametric SEM, considering the following NPSEM [55]
@@
where functions $f_{A},f_{\boldsymbol{X}},f_{Y}$ are no longer restricted to be linear or known. Hence, the wanted estimate is no longer the path coefficients but a mapping from the statistical model to the real line. TMLE for ATE leverages the orthogonal factorization of the likelihood function induced by the NPSEM [44]
@@
We define $Q(A,\boldsymbol{X})=\mathbb{E}(Y|A,\boldsymbol{X})$ and $g(A|\boldsymbol{X})=\mathbb{P}(A|\boldsymbol{X})$, where $Q(A,\boldsymbol{X})$ can be estimated from the observed data and the $g(A|\boldsymbol{X})$ is nuisance parameter that will be used in the subsequent targeting step [44]. For concreteness, we assume the measurement $Y$ is continuous, while other types, such as binary, can also be accommodated with minor modifications [53]. For an unbounded continuous outcome, it is generally recommended to scale $Y$ with the min-max transformation [42, 64]
@@
followed by the inverse transformation after the point and set estimations are completed. The first step in implementing TMLE is to obtain an initial estimate of $\mathbb{E}(Y|A,\boldsymbol{X})$. Since the outcome is bounded in $[0,1]$, one could use logistic regression or machine learning algorithms for estimation, but incorporating cross-validation to avoid overfitting is crucial and generally recommended [45]. Predictions from the fitted or trained model yield $Q_{n}^{0}(A=1,\boldsymbol{X}_{i})$ and $Q_{n}^{0}(A=0,\boldsymbol{X}_{i}),\forall i\in\{1,2,\cdots,n\}$. The targeting step begins with estimating the propensity score, $g(A\mid\boldsymbol{X})$. The clever covariate $H(g,A,\boldsymbol{X})$, derived from the efficient influence function, is defined to be
@@
It has been established that if either $Q^{0}_{n}(A,\boldsymbol{X})$ or $g_{n}(A\mid\boldsymbol{X})$ is consistent, the resulting $\hat{\psi}^{\operatorname{TMLE}}_{\operatorname{ATE}}$ is a consistent estimator for $\psi_{\operatorname{ATE}}$ [64]. To assess the uncertainty of the TMLE, its variance can be estimated using the efficient influence function:
@@
allowing a 95% confidence interval to be constructed as $\hat{\psi}^{\operatorname{TMLE}}_{\operatorname{ATE}}\pm z_{0.975}\hat{\sigma}/\sqrt{n}$ [44]. It has been shown that if both the initial estimate $Q_{n}^{0}$ and the estimated propensity score $g_{n}$ are consistent, then it attains the efficiency [64].
Many tutorials are available for implementing the TMLE for ATE with either a continuous or a binary outcome [42, 53, e.g.,], while the R package tmle also provides a built-in function that can integrate with the SuperLearner [43]. The steps for estimating CATE via TMLE can be done in strata, where the package tmle3 provides a built-in function to achieve this [38].
@@ 3.2 Causal Mediation Analysis
Causal mediation analysis is a statistical framework used to understand how a treatment $A$ affects an outcome $Y$ by decomposing the total causal effect into a direct effect (the part not operating through an intermediate variable) and an indirect effect (the part that operates through a mediator, $M$). This helps disentangle mechanisms in causal pathways, common in fields like epidemiology, psychology, and social sciences. Suppose we have continuous outcome $Y$, a binary treatment or exposure $A\in\mathcal{A}=\{0,1\}$, baseline covariates $\boldsymbol{W}=(W_{1},W_{2},\cdots,W_{p})$ take values in $\mathcal{W}\subseteq\mathbb{R}^{p}$ and a continuous mediator $M$ with possible values $\mathcal{M}\subseteq\mathbb{R}$. We use $Y(a,m)$ to denote the outcome when the treatment $A=a$ and the mediator $M=m$; $M(a)$ represents the mediator with $A=a$. We would make the composition assumption, which states that $Y(a,M(a))=Y(a)$ for $a\in\{0,1\}$ [41]. The ATE in the context of mediation analysis can then be defined similarly
@@
Under the composition assumption, we could decomposite $\psi_{\text{ATE}}=\text{NIE}+\text{NDE}$. To identify the NIE and NDE, we need several assumptions in addition to the standard consistency assumption: (i) sequential ignorability: $A\perp Y(a,m)|\boldsymbol{W}$ and $M\perp Y(a,m)|\boldsymbol{W},A$ for all $a\in\mathcal{A},m\in\mathcal{M}$; (ii) no treatment-mediator confounding: $A\perp M(a)|\boldsymbol{W}$ for all $a\in\mathcal{A}$; and (iii) cross-world independence: $Y(a,m)\perp M(a^{\prime})|\boldsymbol{W}$ for all $a,a^{\prime}\in\mathcal{A},m\in\mathcal{M}$ [41]. The positivity assumption here is extended to be: $\exists\epsilon_{a}>0\text{ such that }\mathbb{P}(A=a\mid\boldsymbol{W}=\boldsymbol{w})>\epsilon_{a}$ $\forall a\in\mathcal{A},\boldsymbol{w}\in\mathcal{W}$ and $\exists\epsilon_{m}>0\text{ such that }\mathbb{P}(M=m\mid A=a,\boldsymbol{W}=\boldsymbol{w})>\epsilon_{m}$ $\forall m\in\mathcal{M},a\in\mathcal{A},\boldsymbol{w}\in\mathcal{W}$. When these assumptions are satisfied, the NDE and NIE can be non-parametrically identified [41, 40]. In particular,

### v2

@@ 1 Introduction
Structural Equation Modeling [bollen-2014, SEM;] is a widely used framework in applied statistics, social sciences, psychology, and related fields [long-2023, szaflarski-2019, hermstad-2010, e.g.,], allowing researchers to model complex interrelationships among observed and latent variables. As a multivariate technique that often incorporates path analysis, SEM also provides an approach to capture mediational mechanisms through direct and indirect effects [curran-2003, gunzler-2013]. In addition to SEM’s adaptability and flexibility, many software implementations are available (e.g., lavaan [rosseel-2012] in R; Stata; and Mplus), which offer further applicability for empirical research.
Numerous research questions in sociology and other fields such as psychology and education could be equivalently answered by estimating the path coefficients within the SEM framework. Under the structural equations and the distributional assumptions on the disturbance terms, parameters are typically estimated by maximizing the likelihood function induced by the model. The subsequent inferences are heavily based on likelihood-based tests or non-parametric approaches such as bootstrap [curran-2003, bollen-2014]. However, as the true data-generating mechanism typically remains unknown, it is not guaranteed that the specified likelihood encodes it correctly [curran-2003]. In the presence of model misspecification, the efficiency and consistency of the estimator and the validity of the corresponding statistical inference are no longer granted, resulting in unreliable estimates and inferences[kaplan-1988, yuan-2003].
@@
Notably, path coefficients and path diagrams (e.g., Figure 1) in SEM can sometimes represent causal effects that are defined under the Neyman–Rubin potential outcomes framework, linking causal inference to parameter estimation [bollen-2013, pearl-2009, e.g.,]. However, the equivalence between path coefficients and causal effects essentially depends on the correct model specification. To address the limitations of linearity and reduce the risk of model misspecification, nonparametric structural equation models [pearl-2009-1, NPSEM;] use nonparametric equations to represent relationships between variables, reformulating path diagrams as Directed Acyclic Graphs (DAGs; Figure 2). Using NPSEM and DAG, causal effects are identified as functionals of the joint distribution rather than path coefficients. Built on NPSEM, Targeted Maximum Likelihood (Loss) Estimation (TMLE), introduced by van der Laan and Rubin [van-der-laan-2011], provides a modern, semiparametric, machine learning-based framework for causal inference that is doubly robust and efficient under certain conditions.
@@
Upon identifying a causal effect from NPSEM, TMLE then involves two steps: an initial data-adaptive estimation followed by an adjustment that targets the estimation equation of the efficient influence function (EIF) [gruber-2009]. TMLE is often coupled with the Super Learner algorithm [van-der-laan-2006]—an ensemble method that combines multiple learners to enhance predictions. Under regularity conditions, TMLE achieves asymptotic efficiency and hence provides a valid uncertainty measurement through the influence function [gruber-2009, van-der-laan-2011], offering computational advantages over bootstrap methods while preserving interpretability.
@@ 2.1 Structural Equation Modeling
[curran-2003] offers a concise review of SEM, from which the following overview is primarily adapted, with minor modifications for notation and clarity. The standard SEM consists of two parts: the measurement submodel and the structural submodel. In general, the structural equation is defined as
@@
Under regularity conditions, the properties of MLE ensure the asymptotic distribution of $\sqrt{n}(\hat{\boldsymbol{\theta}}_{\text{ML}}-\boldsymbol{\theta})$ is multivariate normal with mean $\boldsymbol{0}$ and covariance matrix given by the inverse of the Fisher information matrix, enabling statistical inference such as Wald tests for parameter significance and likelihood ratio tests for model fit. It is possible that the desired estimation involves some nonlinear transformations of the path coefficients. In such a case, the uncertainty is usually captured by the Delta method or non-parametric techniques like bootstrap [bollen-2013].
@@ 2.2.1 An Overview of Targeted Learning
In contrast to parametric SEM, Targeted Maximum Likelihood(Loss) Estimation is a modern, semi-parametric estimation framework developed to provide an efficient and robust estimation of causal effects and other statistical parameters in complex data settings. An NPSEM is generally specified first for describing the causal relations and identifying the causal parameters in terms of the joint distribution. Then, the TMLE involves two main steps: an initial estimation of outcome functions using any suitable method, and a targeting step that updates these estimates to improve the estimation of the parameter of interest. This targeting is done through a fluctuation submodel and solving the estimation equation of the efficient influence function (EIF), which plays a key role in ensuring local efficiency. The resulting TMLE estimator is asymptotically linear and normally distributed under mild regularity conditions, allowing for valid inference via standard error estimation and confidence intervals. Following [gruber-2010], we assume that a semi-parametric statistical model has been identified $\mathcal{M}$ with a true but unknown distribution $F_{0}\in\mathcal{M}$. The interested parameter is defined as $\Psi(F_{0}),\Psi(\cdot):\mathcal{M}\rightarrow\mathbb{R}$, and with $O_{1},O_{2},\cdots,O_{n}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}F_{0}$ observations from $F_{0}$. We further need that $\Psi(F_{0})=\Psi(Q_{0})$ depends on only $Q_{0}=Q(F_{0})$, a part of $F_{0}$, and $g_{0}$ is a nuisance parameter from some orthogonal factorization.
@@
which also yields a valid confidence set. This framework has been adapted to many estimands that admit the asymptotic linear estimation (e.g. [gruber-2010, hejazi-2022, stitelman-2012]).
@@ 2.2.2 The Super Learner
As mentioned, the efficiency and consistency of TMLE depend heavily on the initial estimates of $Q_{n}^{0}$ and $g_{n}$. Though machine learning algorithms could be applied instead of the parametric regression, we by no means should rely on a single algorithm, as none is optimal for all datasets. It is worth mentioning that the estimation of $Q_{n}^{0}$ and $g_{n}$ can be done non-parametrically via the Super Learner [van-der-laan-2007], an ensemble learning technique that integrates modern machine learning algorithms. Consider a dataset $\{(X_{i},Y_{i})\}_{i=1}^{n}$, where $X_{i}\in\mathbb{R}^{p}$ are features and $Y_{i}\in\mathbb{R}$ are outcomes, jointly drawn from an unknown distribution $F_{0}\in\mathcal{M}$. Recall that the initial estimates in TMLE are equivalent to constructing a function $f(X)$ that minimizes the expected risk
@@
Figure 3: Illustration of the Super Learner algorithm flow. Adapted from [van-der-laan-2007].
Numerous modern machine learning algorithms are included in the current SuperLearner implementation in R, including Random Forest, Support Vector Machine, eXtreme Gradient Boosting, and Artificial Neural Network, which enhances its versatility and applicability. Notably, the choices of base learners and the meta learner play a key role in the final estimation, and have to be considered based on the specific tasks involved [phillips-2023].
@@ 3 SEM and TMLE approach to causal inference
When considering SEM for causal inference, the commonly considered parameters are the Average Treatment Effect (ATE), Conditional Average Treatment Effect (CATE), and mediational effects [gunzler-2013, bollen-2013]. In this section, we consider these causal effects and show that both SEM and TMLE are applicable for estimation.
@@ 3.1 On Average Treatment Effects
Three essential causal assumptions for efficient identification and estimation of the causal effect from accessible data are (i) Stable Unit Treatment Value Assumption (SUTVA) or the consistency assumption: $Y=A\cdot Y(1)+(1-A)\cdot Y(0)$; (ii) no unmeasured confounders: $\mathbb{P}(A=1|Y(1),Y(0),\boldsymbol{X})=\mathbb{P}(A=1|\boldsymbol{X})$; and (iii) positivity: For all $a\in\mathcal{A}$, $\exists\epsilon>0\text{ such that }\mathbb{P}(A=a|\boldsymbol{X}=\boldsymbol{x})>\epsilon$ for almost all $\boldsymbol{x}\in\mathcal{X}$, without which the ATE might not be identifiable [rosenbaum-1983].
@@ 3.1.1 The SEM approach to ATE
Suppose that the data-generating mechanism is encoded by the structural linear causal models [pearl-2009]
@@
where $\ell$ denotes the log-likelihood function with $\phi(\mu,\tau^{2})$ denotes the normal density with mean $\mu$ and variance $\tau^{2}$, and $O$ represents the observed data, or as part of a generalized SEM with the logit link. It is unbiased, consistent, efficient, and asymptotically normal under correct model specification [wasserman-2013]. The statistical inference regarding the ATE $H_{0}:\gamma=0$, in a SEM context, is usually done via the Wald test and the likelihood ratio test.
@@
If $X_{k}=x_{k}$ is constant in $B$, the CATE simplifies to $\gamma+\tau x_{k}$. Estimation follows similarly, with asymptotic normality preserved under linear transformations, and the inference remains valid up to a minor modification [wasserman-2013].
@@ 3.1.2 The targeted learning approach to ATE
However, the identification of causal effects as path coefficients and the consistency of the estimators rely on correct functional form specification, and misspecification can introduce bias[curran-2003, yuan-2003]. In contrast to parametric SEM, considering the following NPSEM [pearl-2009]
@@
where functions $f_{A},f_{\boldsymbol{X}},f_{Y}$ are no longer restricted to be linear or known. Hence, the wanted estimate is no longer the path coefficients but a mapping from the statistical model to the real line. TMLE for ATE leverages the orthogonal factorization of the likelihood function induced by the NPSEM [gruber-2009]
@@
We define $Q(A,\boldsymbol{X})=\mathbb{E}(Y|A,\boldsymbol{X})$ and $g(A|\boldsymbol{X})=\mathbb{P}(A|\boldsymbol{X})$, where $Q(A,\boldsymbol{X})$ can be estimated from the observed data and the $g(A|\boldsymbol{X})$ is nuisance parameter that will be used in the subsequent targeting step [gruber-2009]. For concreteness, we assume the measurement $Y$ is continuous, while other types, such as binary, can also be accommodated with minor modifications [luquefernandez-2018]. For an unbounded continuous outcome, it is generally recommended to scale $Y$ with the min-max transformation [frank-2023, van-der-laan-2011]
@@
followed by the inverse transformation after the point and set estimations are completed. The first step in implementing TMLE is to obtain an initial estimate of $\mathbb{E}(Y|A,\boldsymbol{X})$. Since the outcome is bounded in $[0,1]$, one could use logistic regression or machine learning algorithms for estimation, but incorporating cross-validation to avoid overfitting is crucial and generally recommended [gruber-2010]. Predictions from the fitted or trained model yield $Q_{n}^{0}(A=1,\boldsymbol{X}_{i})$ and $Q_{n}^{0}(A=0,\boldsymbol{X}_{i}),\forall i\in\{1,2,\cdots,n\}$. The targeting step begins with estimating the propensity score, $g(A\mid\boldsymbol{X})$. The clever covariate $H(g,A,\boldsymbol{X})$, derived from the efficient influence function, is defined to be
@@
It has been established that if either $Q^{0}_{n}(A,\boldsymbol{X})$ or $g_{n}(A\mid\boldsymbol{X})$ is consistent, the resulting $\hat{\psi}^{\operatorname{TMLE}}_{\operatorname{ATE}}$ is a consistent estimator for $\psi_{\operatorname{ATE}}$ [van-der-laan-2011]. To assess the uncertainty of the TMLE, its variance can be estimated using the efficient influence function:
@@
allowing a 95% confidence interval to be constructed as $\hat{\psi}^{\operatorname{TMLE}}_{\operatorname{ATE}}\pm z_{0.975}\hat{\sigma}/\sqrt{n}$ [gruber-2009]. It has been shown that if both the initial estimate $Q_{n}^{0}$ and the estimated propensity score $g_{n}$ are consistent, then it attains the efficiency [van-der-laan-2011].
Many tutorials are available for implementing the TMLE for ATE with either a continuous or a binary outcome [frank-2023, luquefernandez-2018, e.g.,], while the R package tmle also provides a built-in function that can integrate with the SuperLearner [gruber-2012]. The steps for estimating CATE via TMLE can be done in strata, where the package tmle3 provides a built-in function to achieve this [Jeremy-2021].
@@ 3.2 Causal Mediation Analysis
Causal mediation analysis is a statistical framework used to understand how a treatment $A$ affects an outcome $Y$ by decomposing the total causal effect into a direct effect (the part not operating through an intermediate variable) and an indirect effect (the part that operates through a mediator, $M$). This helps disentangle mechanisms in causal pathways, common in fields like epidemiology, psychology, and social sciences. Suppose we have continuous outcome $Y$, a binary treatment or exposure $A\in\mathcal{A}=\{0,1\}$, baseline covariates $\boldsymbol{W}=(W_{1},W_{2},\cdots,W_{p})$ take values in $\mathcal{W}\subseteq\mathbb{R}^{p}$ and a continuous mediator $M$ with possible values $\mathcal{M}\subseteq\mathbb{R}$. We use $Y(a,m)$ to denote the outcome when the treatment $A=a$ and the mediator $M=m$; $M(a)$ represents the mediator with $A=a$. We would make the composition assumption, which states that $Y(a,M(a))=Y(a)$ for $a\in\{0,1\}$ [ding-2024]. The ATE in the context of mediation analysis can then be defined similarly
@@
Under the composition assumption, we could decomposite $\psi_{\text{ATE}}=\text{NIE}+\text{NDE}$. To identify the NIE and NDE, we need several assumptions in addition to the standard consistency assumption: (i) sequential ignorability: $A\perp Y(a,m)|\boldsymbol{W}$ and $M\perp Y(a,m)|\boldsymbol{W},A$ for all $a\in\mathcal{A},m\in\mathcal{M}$; (ii) no treatment-mediator confounding: $A\perp M(a)|\boldsymbol{W}$ for all $a\in\mathcal{A}$; and (iii) cross-world independence: $Y(a,m)\perp M(a^{\prime})|\boldsymbol{W}$ for all $a,a^{\prime}\in\mathcal{A},m\in\mathcal{M}$ [ding-2024]. The positivity assumption here is extended to be: $\exists\epsilon_{a}>0\text{ such that }\mathbb{P}(A=a\mid\boldsymbol{W}=\boldsymbol{w})>\epsilon_{a}$ $\forall a\in\mathcal{A},\boldsymbol{w}\in\mathcal{W}$ and $\exists\epsilon_{m}>0\text{ such that }\mathbb{P}(M=m\mid A=a,\boldsymbol{W}=\boldsymbol{w})>\epsilon_{m}$ $\forall m\in\mathcal{M},a\in\mathcal{A},\boldsymbol{w}\in\mathcal{W}$. When these assumptions are satisfied, the NDE and NIE can be non-parametrically identified [ding-2024, de-stavola-2014]. In particular,